OpenStudy (anonymous):

A bag of 125 stickers contains 24% azure, 28% pink, 18% lime, and 30% amber stickers. Write a proportion that can be used to find L, the number of lime stickers in the bag.

OpenStudy (anonymous):

Ok, my computer's freezing up, so idk if youre sayig anything!

OpenStudy (lgbasallote):

im not saying anything...yet....

OpenStudy (anonymous):

Okiedokie, sorry my computer was messing up

OpenStudy (lgbasallote):

i know how to get the number of lime stickers...but i dont know how to type it as proportion...

OpenStudy (anonymous):

Oh, well I was think something like 18/125=l/100 but idk.

OpenStudy (lgbasallote):

hmm there may be a way....do you know how to find 18% of 125?

OpenStudy (anonymous):

125 times .18?

OpenStudy (lgbasallote):

and how did you get 0.18?

OpenStudy (anonymous):

Well to multiply a percent you move the decimal twice to the left...

OpenStudy (anonymous):

So 125 times .18=22.5

OpenStudy (lgbasallote):

or...divide by 100 so... \[\Large \frac{8}{100} \times 125 = L\] agree?

OpenStudy (lgbasallote):

that should be 18 not 8

OpenStudy (anonymous):

So how would I make that a proportion?

OpenStudy (lgbasallote):

\[\Large \frac{18}{100} = \frac{L}{125}\] it looks like a proportion now! dont forget to simplify...if you want you can change it to the form a:b = c:d thingy

OpenStudy (anonymous):

Okiedokie, thanks again! :D